# LDI: Localized Data Imputation

### v1

@@ IV Proposed Method
LLMs have shown strong performance on a range of table reasoning tasks, making them promising candidates for data imputation. However, they often struggle with scalability to large tables [43, 44] and lack interpretability in their imputation decisions. To address these issues, our method, LDI, selectively identifies the most relevant attributes and tuples to focus the model’s attention and improve explainability. In the first phase (§$\S$§IV-A), the method captures and retains only the attributes that have an underlying connection with $A_{T}$ and can contribute to the imputation process. The second phase (§$\S$§IV-B) implements a tuple-level filtering mechanism to retain a diverse and relevant set of records. Finally, in the third phase (§$\S$§IV-C), the refined attributes and tuples are fed into the LLM, leveraging it as a function for imputing missing values. An overview of this process is demonstrated in Figure 2.
Figure 2: Overview of the LDI framework
IV-A Attribute Selection
@@
To identify attributes most relevant to the imputation target, LDI performs attribute selection in three phases: (1) selecting a representative subset of tuples, (2) detecting group-specific patterns in candidate columns, and (3) evaluating approximate dependency relationships. This section introduces the relaxed dependency criterion we use and outlines how it guides attribute selection.
@@
Traditional functional dependencies are commonly used to model relationships between attributes. However, they tend to be too strict for real-world tables that contain noise and long textual values. To better handle such cases, we introduce a relaxed dependency criterion tailored to noisy, text-rich data.
@@
We use this criterion to identify a subset of candidate attributes $A_{C}$ that exhibit a $(p,q)$-approximate dependency with the target attribute $A_{T}$ with the caveat that such dependencies, like traditional functional dependencies, cannot be definitively inferred from a database instance, and one can only verify that a possible dependency is not violated. Verifying these dependencies can be computationally expensive, especially as the number of attributes and tuples grows.
To reduce computational overhead, LDI operates on a sampled subset of tuples. Naive random sampling is inadequate as it overrepresents frequent values in $A_{T}$ and underrepresents rare ones, which can obscure group-level patterns. Since LDI relies on identifying such patterns, it is essential to preserve the diversity of $A_{T}$ in the sample. We address this with group sampling, which ensures that values across different groups are proportionally represented.
IV-A1 Group Sampling
To maintain the diversity of the target attribute $A_{T}$, we sample data in a structured way. We first partition the table into groups based on distinct values in $A_{T}$. From these, we randomly select $m$ groups that each contain at least $n$ rows. Within each selected group, we randomly sample $n$ rows, yielding a total of $m\times n$ records.
The parameters $m$ (number of groups) and $n$ (number of tuples per group) control the breadth and depth of the sample. Larger values of $m$ help capture a wider range of distinct $A_{T}$ values, while higher $n$ supports the detection of more subtle or noisy within-group patterns. Smaller values may suffice for more uniform datasets. The goal is to retain meaningful group-level signals while keeping the overall sample size manageable, and this stratified approach ensures that less frequent values of $A_{T}$ are not overlooked, supporting more reliable dependency detection in later stages.
In the next step, LDI uses this sampled data to identify signals in candidate attributes that point to a dependency on the target attribute.
IV-A2 Pattern Detection
@@
To assess whether a candidate attribute $A_{C}$ exhibits a dependency on the target attribute $A_{T}$, we analyze its values for group-specific patterns. We use the Longest Common Substring (LCS) to detect recurring character sequences—such as prefixes, category tags, or embedded keywords—that frequently appear within groups sharing the same $A_{T}$ value. These patterns may not hold globally, but they provide localized signals that are informative for imputation.
Within each group (i.e., a set of tuples sharing the same $A_{T}$ value), we examine the values of $A_{C}$ to extract all substrings. A substring is retained if it appears in at least a $q$ fraction of the group’s tuples. This threshold filters out infrequent or incidental patterns and can be tuned to balance noise tolerance with pattern sensitivity.
@@
Each frequent substring in $A_{C}$ must be uniquely associated with a single value in the $A_{T}$ column; otherwise, it violates the functional dependency condition. To enforce this constraint, we discard any substring that appears as frequent in more than one group.
@@
Additionally, we remove redundant substrings that are fully contained in longer, more informative ones within each group. Since frequent substrings imply that all their subparts are also frequent, retaining them only increases clutter. This cleanup step is optional, as it does not affect the correctness of dependency detection, but it simplifies interpretation and highlights the most distinctive patterns when explaining why an attribute was selected.
Example 2. Consider a group where the $A_{C}$ values are {‘aab’, ‘aa’, ‘ab’}. The algorithm extracts all substrings and counts their frequency across the three tuples: ‘a’ appears in all 3 tuples. ‘aa’, ‘ab’, and ‘b’ each appear in 2 tuples. ‘aab’ appears in 1 tuple. With a threshold $q=0.6$, substrings that appear in at least 2 of the 3 tuples are retained: {‘a’, ‘b’, ‘aa’, ‘ab’}. Next, substrings common across other groups are removed. Suppose ‘a’ appears frequently in other groups and is excluded: {‘aa’, ‘ab’, ‘b’}. Then, shorter substrings that are part of longer retained ones are discarded. Since ‘b’ is a substring of ‘ab’, it is removed. The final patterns for this group are: {‘aa’, ‘ab’}.
@@
The LCS method is particularly effective in uncovering non-trivial patterns, even in the presence of noises such as typos or formatting inconsistencies. Its robustness stems from two key properties: (1) If a pattern is absent in some cells of a group in $A_{C}$, the LCS remains intact for others. By adjusting the parameter $q$, one can reduce the required number of matching cells while still detecting meaningful dependencies. (2) Inconsistent formatting produces shorter but still relevant LCS results. For example, if grouped cells contain ‘+1702-’, ‘702/’, and ‘702-’, the extracted substring ‘702’ (the area code) still reveals the underlying pattern (see Figure 3).
IV-A3 Dependency Evaluation
@@
In the final step of the attribute selection phase, we evaluate whether there is a $(p,q)$-approximate dependency between $A_{C}$ and $A_{T}$. In this step, the algorithm checks how many groups have at least one unique pattern (i.e., a substring, regardless of its length, that frequently appears in that group but is not commonly found in other groups). If such patterns are found in at least a fraction $p$ of all groups, the attribute $A_{T}$ is considered dependent on $A_{C}$. This condition ensures that the dependency between the two attributes is supported across multiple groups, rather than being influenced by random or rare patterns.
Example 3. Suppose we are evaluating the dependency between a candidate attribute $A_{C}$ and the target attribute $A_{T}$, which contains 5 distinct values (i.e., 5 groups). After applying the pattern detection step, where only group-specific substrings are retained, we obtain the following LCSs: {‘aa’, ‘ab’}, {‘cc’, ‘ac’}, {‘d’}, { } [^2] This is a group where no common substring with sufficient frequency is retained., {‘ee’}. Each substring shown is unique to its respective group, as ensured by the previous step. To evaluate the dependency between $A_{C}$ and $A_{T}$, we count how many groups contain at least one retained pattern. In this example, Groups 1, 2, 3, and 5 satisfy this condition, yielding 4 out of 5 supporting groups. For any value of $p\leq 0.8$, the algorithm concludes that a meaningful dependency exists between $A_{C}$ and $A_{T}$.
This procedure is repeated for all candidate attributes, and those with detected dependencies are passed to the next phase. Algorithm 1 summarizes the key steps of the attribute selection.
Input: Dataset $\mathcal{D}$, Attribute $A_{T}$
Output: Set of selected attributes
1
@@
3 $\mathcal{D}^{\prime}\leftarrow\text{Sample}(\mathcal{D})$;
4 Group $\mathcal{D}^{\prime}$ by unique values in $A_{T}$;
5
6foreach $A_{C}\in\mathcal{R}\setminus A_{T}$ do
7       foreach group $g$ in $\mathcal{D}^{\prime}$ do
8             Generate all substrings;
9             Count occurrences of each substring;
10             Retain substrings appear in at least $q$ fraction of records;
11
12      Identify unique substrings for each group $g$;
13       if At least $p$ fraction of groups have unique substrings then
[…]
15
16
17return selected_attributes;
Algorithm 1 Attribute Selection (Simplified)
IV-B Tuple Selection
@@
In the second phase of our proposed method, we are looking for a small set of tuples as examples to guide the LLM. The selection process follows three key criteria: (1) the tuples must have complete values in $A_{T}$, (2) they should be similar to the tuple with the missing value to provide relevant context, and (3) be diverse in their $A_{T}$ values to reduce the risk of bias.
The dataset is first filtered to include only tuples with complete values in $A_{T}$, as they are needed to serve as valid examples. To ensure focus on informative patterns and avoid distractions from irrelevant fields, we then restrict the similarity computation to the selected attributes identified in the first phase and take the mean similarity across selected attributes. For a tuple $t_{i}$ with $t_{i}[A_{T}]=\varnothing$, its similarity to a complete tuple $t_{j}$ is defined as:
$\text{similarity}(t_{i},t_{j})=\frac{1}{|S|}\sum_{A\in S}\frac{|\text{LCS}(t_{i}[A],t_{j}[A])|}{\max(|t_{i}[A]|,|t_{j}[A]|)}$ (1)
where $S$ is the set of relevant attributes selected in the first phase, and $\text{LCS}(t_{i}[A],t_{j}[A])$ computes the longest common substring between the values of attribute $A$ in tuples $t_{i}$ and $t_{j}$.
Finally, we select the top $k$ tuples with distinct values in $A_{T}$ from this ranking to balance similarity with diversity among the chosen examples and avoid over-representation of any one category. These $k$ examples form the input context and are passed to the LLM in the next phase, where they are used to impute the missing value in $t_{i}$ through few-shot prompting.
@@
Example 4. Suppose the relevant attributes selected in the first phase are $A_{1}$ and $A_{2}$, and we want to impute the missing value in $A_{T}$ for the tuple $t_{i}$.Table I presents the tuples used for imputation.
$A_{1}$ $A_{2}$ $A_{T}$
$t_{p}$ ab defg X
$t_{q}$ bc fg Y
$t_{r}$ cab ef X
$t_{i}$ abc def N/A
TABLE I: Example of selecting candidate tuples
- Similarity between $t_{i}$ and $t_{p}$:
LCS(‘abc’, ‘ab’) = ‘ab’ $\rightarrow$ length = 2, max length = 3 $\rightarrow$ 2/3
LCS(‘def’, ‘defg’) = ‘def’ $\rightarrow$ length = 3, max length = 4 $\rightarrow$ 3/4
Average similarity = (2/3 + 3/4) / 2 $\approx$ 0.708
- Similarity between $t_{i}$ and $t_{q}$:
LCS(‘abc’, ‘bc’) = ‘bc’ $\rightarrow$ length = 2, max length = 3 $\rightarrow$ 2/3
LCS(‘def’, ‘fg’) = ‘f’ $\rightarrow$ length = 1, max length = 3 $\rightarrow$ 1/3
Average similarity = (2/3 + 1/3) / 2 = 0.500
- Similarity between $t_{i}$ and $t_{r}$:
LCS(‘abc’, ‘cab’) = ‘ab’ $\rightarrow$ length = 2, max length = 3 $\rightarrow$ 2/3
LCS(‘def’, ‘ef’) = ‘ef’ $\rightarrow$ length = 2, max length = 3 $\rightarrow$ 2/3
Average similarity = (2/3 + 2/3) / 2 = 0.667
- Ranking based on similarity:
  1. 1.
    $t_{p}$: 0.708 ($A_{T}$ = ‘X’)
  2. 2.
    $t_{r}$: 0.667 ($A_{T}$ = ‘X’)
  3. 3.
    $t_{q}$: 0.500 ($A_{T}$ = ‘Y’)
When $k=1$, the method selects the tuple $t_{p}$, which has the highest similarity to $t_{i}$. When $k=2$, it selects $t_{p}$ and $t_{q}$ to preserve similarity and reduce bias by covering different values of $A_{T}$.
IV-C Data Imputation Using LLM
@@
In the third phase of our proposed method, we leverage LLMs to impute missing values. This phase employs few-shot learning by providing the LLM with a small set of example tuples that include the relevant attributes identified in § IV-A and retrieved similar tuples from § IV-B. The LLM then infers the missing value by recognizing patterns across these examples.
@@
To prepare the input, we adopt a standard serialization format from prior work [13], where each tuple is expressed as a sequence of key-value pairs. The input prompt consists of a task description, a list of examples, and a query row with a missing value:
@@
Once the input is constructed, the LLM uses the examples to infer the missing value for the target attribute $A_{T}$. This allows the model to generalize from limited data and detect complex relationships between attributes.
@@
Example 1 (Motivating Example) continued: Our proposed method first groups the data by the values in the City column and selects $m$ groups, each with $n$ records, to generate a representative sample (see Figure 3). It then iterates through all candidate columns, applying the pattern detection method to identify repeated substrings that may serve as group-specific signals. In this example, both the Phone and Reviews columns exhibit recurring patterns within groups. However, during the dependency evaluation step, only the Phone column is selected. Although the Reviews column contains the substring “[(’Rated ” across several groups, this pattern is not unique to any specific city, and thus does not meet the criteria for group-level dependency. In contrast, the Phone column shows city-specific area codes that strongly correlate with the target values, and is therefore retained for imputation. Importantly, the method can handle formatting inconsistencies in the data. For example, phone numbers may include different prefixes (e.g., +1) or use varied separators. Despite such variations, our LCS-based approach is able to capture the consistent patterns, such as the core area code, across the grouped entries.
In the second phase, the method uses the selected attribute(s), in this case, Phone, to identify tuples similar to the one with the missing value. It applies a similarity search and selects $k$ diverse examples that are most similar to the incomplete tuple while having distinct values in the target column. This ensures that the chosen examples are both relevant and diverse, providing a balanced context for the final prediction.
In the third phase, these selected tuples are serialized into key-value formats and provided to an LLM in a few-shot prompting setup. The LLM is asked to predict the missing city value based on the patterns observed in the provided examples. By focusing only on relevant attributes and carefully selected examples, the model is guided toward accurate and more explainable imputations. In this example, the LLM would rely solely on the signals in the Phone column, making the basis of the imputed value both precise and traceable. This illustrates how LDI can effectively reduce noise, focus on informative data, and produce traceable results even in the absence of predefined relationships in the dataset.
Figure 3: Example of data imputation (after applying our approach)
V Performance Analysis

### v2

@@ 4. Proposed Method
While LLMs’ strong performance on a range of table reasoning tasks makes them promising candidates for data imputation, they often struggle with scalability to large tables (Chen, 2023; Ji et al., 2023) and lack interpretability in their imputation decisions. To address these issues, LDI, selectively identifies the most relevant attributes and tuples, focusing the model’s attention and improving explainability. In the first phase ($\S$4.1), LDI retains only the attributes that are likely related to the target attribute $A_{T}$ and informative for imputation. The second phase ($\S$4.2) filters tuples to produce a diverse and relevant set of records. Finally, in the third phase ($\S$4.3), the selected attributes and tuples are fed into the LLM, which performs the imputation. An overview of this process is demonstrated in Figure 2.
Figure 2. Overview of the LDI framework
4.1. Attribute Selection
@@
To identify attributes most relevant to the imputation target, LDI performs attribute selection in two main phases: (1) detecting group-specific patterns in candidate columns, and (2) evaluating approximate dependency relationships.
@@
While traditional functional dependencies can model relationships between attributes, they are often too strict for real-world tables that contain noise and long textual values. To better handle such cases, we introduce a relaxed dependency criterion tailored to noisy, text-rich data.
@@
We use this criterion to identify a subset of candidate attributes $A_{C}$ that exhibit a $(p,q)$-approximate dependency with the target attribute $A_{T}$ with the caveat that such dependencies, like traditional functional dependencies, cannot be definitively inferred from a database instance, and one can only verify that a possible dependency is not violated.
4.1.1. Pattern-Based Dependency Detection
@@
To assess whether a candidate attribute $A_{C}$ exhibits a dependency on the target attribute $A_{T}$, LDI analyzes its values for group-specific patterns. We use the Longest Common Substring (LCS) to detect recurring character sequences—such as prefixes, category tags, or embedded keywords—that frequently appear within groups sharing the same $A_{T}$ value. These patterns provide localized signals that are informative for imputation, even if they do not hold globally.
@@
Within each group (i.e., tuples sharing the same $A_{T}$ value), we extract all substrings from $A_{C}$. A substring is retained if it appears in at least a $q$ fraction of the group’s tuples, therby filtering out infrequent or incidental patterns while balancing robustness to noise with sensitivity to meaningful regularities. Each frequent substring in $A_{C}$ must be uniquely associated with a single value in the $A_{T}$ column; otherwise, it violates the functional dependency condition. To enforce this constraint, we discard any substring that appears as frequent in more than one group.
@@
Additionally, we remove redundant substrings that are fully contained in longer, more informative ones within each group. Since every frequent substring implies that all of its subparts are also frequent, retaining them only increases clutter. While this cleanup step is optional and does not affect the correctness, it simplifies interpretation and highlights the most distinctive patterns for explainability.
For example, consider the Phone attribute in our running example (Figure 1). Las Vegas phone numbers (e.g., 702/876-4190, 702/731-4036, +1702-385-7111) share a common prefix pattern, whereas New York numbers (212/333-7788, 212/431-0021, 212/688-5447) follow a different pattern. With a threshold $q=0.6$, substrings such as $\{\mbox{‘702/’},\mbox{‘702’},\mbox{‘02/’},\mbox{‘70’},\ldots\}$ are retained for Las Vegas, and $\{\mbox{‘212/’},\mbox{‘212’},\mbox{‘12/’},\mbox{‘21’},\ldots\}$ for New York. Substrings that appear across groups or are fully contained within longer patterns are discarded, leaving only the most distinctive patterns (e.g., ‘702/’ for Las Vegas and ‘212/’ for New York).
@@
The LCS method is particularly effective in uncovering non-trivial patterns, even in the presence of noises such as typos or formatting inconsistencies. Its robustness stems from two key properties: (1) If a pattern is missing in some cells of a group, the LCS across the remaining cells is preserved. Adjusting the parameter $q$ makes it possible to relax the matching requirement while maintaining the detection of meaningful dependencies. (2) Inconsistent formatting produces shorter but still relevant LCS results. For example, variations such as ‘+1780 ’, ‘780/’, and ‘780-’, strill reveal the underlying pattern ‘780’.
4.1.2. Dependency Evaluation
@@
In the second step of the attribute selection phase, we evaluate whether a $(p,q)$-approximate dependency exists between $A_{C}$ and $A_{T}$. Specifically, the algorithm counts how many groups have at least one unique pattern—i.e., a substring that frequently appears within that group but is uncommon across others. If such patterns are found in at least a fraction $p$ of all groups, we consider $A_{T}$ to be approximately dependent on $A_{C}$. This criterion ensures that the inferred dependency is supported by consistent evidence across multiple groups, rather than arising from isolated or random occurrences.
@@
This process is applied to all candidate attributes, and those satisfying the $(p,q)$-dependency condition are retained for the next phase.
4.1.3. Scalability through Stratified Group Sampling
Although our method can efficiently detect dependencies on large datasets (see $\S$5.1 and $\S$6.6), datasets with hundreds of thousands or millions of records may still require additional measures to ensure scalability. To this end, we introduce an optional group sampling step.
While LDI can operate on a random sample of tuples, naive random sampling is inadequate as it overrepresents frequent values in $A_{T}$ and underrepresents rare ones, which can obscure group-level patterns. Since LDI relies on identifying such patterns, it is essential to preserve the diversity of $A_{T}$ in the sample.
Group Sampling. To maintain diversity in the target attribute $A_{T}$, we partition the table by distinct values of $A_{T}$ and then sample in a structured way. We randomly select $m$ groups that each contain at least $n$ rows, and within each group, we randomly draw $n$ rows, producing a sample of $m\times n$ tuples.
When the dataset lacks sufficient tuples for a full $m\times n$ sample, such as when some values of $A_{T}$ are infrequent, we include as many complete groups as possible, prioritizing those with larger sizes. Specifically, we first select all groups with at least $n$ rows, then add those with n1n-1 rows, continuing until no further groups can be added without exceeding the sample limit.
The parameters $m$ (number of groups) and $n$ (tuples per group) control the breadth and depth of the sample. Increasing $m$ covers more distinct $A_{T}$ values, while increasing $n$ captures finer-grained or noisier within-group patterns. Smaller values of $(m,n)$ may suffice for more homogeneous datasets.
This stratified approach preserves group-level diversity while keeping the sample size manageable, enabling LDI to scale effectively without sacrificing the reliability of dependency detection.
4.2. Tuple Selection
@@
In the second phase of our proposed method, we aim to identify a small set of tuples to serve as examples for guiding the LLM during imputation. The selection process follows three key criteria: (1) tuples must have complete values in $A_{T}$, (2) they should be similar to the tuple with the missing value to provide relevant context, and (3) they should be diverse in their $A_{T}$ values to reduce the risk of bias.
We first filter the dataset to retain only tuples with complete values in $A_{T}$, as they are needed to serve as valid examples [^2] Tuples may have missing values in other attributes since full completeness is not required, but they are less likely to be similar to the selected tuple for imputation.. To avoid distractions from irrelevant attributes, similarity is computed over the subset of relevant attributes identified in the previous phase. For a tuple $t_{i}$ with a missing $A_{T}$ value, its similarity to a complete tuple $t_{j}$ is measured using the normalized length of their LCS across the selected attributes $S$:
(1) $\text{similarity}(t_{i},t_{j})=\frac{1}{|S|}\sum_{A\in S}\frac{|\text{LCS}(t_{i}[A],t_{j}[A])|}{\max(|t_{i}[A]|,|t_{j}[A]|)}$
Finally, we select the top $k$ tuples with distinct values in $A_{T}$, balancing similarity with diversity to ensure the examples are both relevant and representative. These $k$ examples are then passed to the LLM for few-shot imputation.
@@
Suppose the relevant attribute selected in the first phase for our running example is Phone, and we want to impute the missing City for tuple $t_{i}$ corresponding to Moongate in Figure 1. The following table presents the selected subset:
Tuple Phone City
$t_{1}$ 212/333-7788 New York
$t_{2}$ 702/876-4190 Las Vegas
$t_{3}$ 212/431-0021 New York
$t_{4}$ 702/731-4036 Las Vegas
$t_{5}$ +1702-385-7111 Las Vegas
$t_{6}$ 212/688-5447 New York
$t_{i}$ 702-791-7352 N/A
We compute the pairwise similarity of $t_{i}$ with each complete tuple based on their Phone values as follows:
Pair LCS Length Max Length Similarity
$t_{i}$, $t_{1}$ -7 2 12 0.167
$t_{i}$, $t_{2}$ 702 3 12 0.250
$t_{i}$, $t_{3}$ 02 2 12 0.167
$t_{i}$, $t_{4}$ 702 3 12 0.250
$t_{i}$, $t_{5}$ 702- 4 14 0.286
$t_{i}$, $t_{6}$ 7 1 12 0.083
Ranking the tuples by similarity gives:
  1. (1)
    $t_{5}$: 0.286 (City = Las Vegas)
  2. (2)
    $t_{2}$: 0.250 (City = Las Vegas)
  3. (3)
    $t_{4}$: 0.250 (City = Las Vegas)
  4. (4)
    $t_{1}$: 0.167 (City = New York)
  5. (5)
    $t_{3}$: 0.167 (City = New York)
  6. (6)
    $t_{6}$: 0.083 (City = New York)
When $k=1$, the most similar tuple $t_{5}$ is selected. When $k=2$, the algorithm chooses $t_{5}$ and $t_{1}$ to preserve similarity while covering different City values.
4.3. Data Imputation Using LLM
@@
In the third phase of LDI, we leverage LLMs to impute missing values using few-shot learning. The model is provided with a small set of example tuples that include the relevant attributes identified in § 4.1 and retrieved similar tuples from § 4.2. The LLM then infers the missing value by recognizing patterns across these examples.
@@
Following prior work (Narayan et al., 2022), each tuple is serialized as a sequence of key-value pairs. The input prompt consists of a task description, example tuples, and a query tuple with a missing value:
@@
The LLM leverages the provided examples to predict the missing value for the target attribute $A_{T}$, enabling it to capture complex dependencies, handle noisy data, and infer plausible values even for patterns not seen in the examples.
@@
Running Example Recap. Suppose we aim to impute the missing City value for the tuple $t_{i}$ corresponding to Moongate in Figure 1. The method first groups the data by City and samples $m$ groups with $n$ tuples each (Figure 3). Pattern detection identifies recurring substrings in candidate columns. Here, both Phone and Reviews show repeated substrings, but only Phone satisfies the group-level dependency criteria: area codes are unique to cities, whereas patterns in Reviews (e.g., “[(’Rated ”) appear across multiple cities and are discarded. Next, similarity search identifies $k$ tuples most relevant to $t_{i}$ based on the selected attribute (Phone) while ensuring diversity in City values. These tuples are serialized into key-value format and provided to the LLM in a few-shot setup. The model then imputes the missing city by relying solely on informative signals in the Phone column, producing predictions that are both accurate and traceable. This workflow demonstrates how LDI reduces noise, focuses on informative data, and leverages LLM reasoning to produce explainable imputations, even without predefined dependencies in the dataset.
Figure 3. Example of data imputation (after applying our approach)
5. Performance Analysis